OpenStudy (anonymous):

Simply the expotential expression 9²

ganeshie8 (ganeshie8):

\(9^2 = 9 * 9 = ?\)

OpenStudy (anonymous):

81?

ganeshie8 (ganeshie8):

yep !! there 2 is called the exponent. it works like this : \(9^3 = 9 * 9 * 9 \) \(9^4 = 9 * 9 * 9 * 9\) \(9^5 = 9 * 9 * 9 * 9 * 9\)

OpenStudy (anonymous):

Okay, thank you for your help :)

ganeshie8 (ganeshie8):

yw.. ^_^